# Instant Runoff Voting and the Reinforcement Paradox

### v1

@@ 5.2. Results
Table 10 also displays information for three large $\alpha$ values which are not shown in Figure 2. Recall that as $\alpha\rightarrow\infty$ the distribution resembles the IC model, in which each voter independently chooses one of the six possible rankings at random. For comparison, we also conducted simulations using the IC model with $V=1001$, which returned an estimated probability of 1.0 for the RU, PL or both RU and PL paradoxes occurring. Intuitively, this is clear since when every voter chooses a ranking at random, $X_{1}\approx X_{2}\approx Y_{1}\approx Y_{2}\approx Z_{1}\approx Z_{2}$. Thus, in the limit, Condition (2) and the plurality loser inequalities of Proposition 1 will always be satisfied. It is unsurprising, therefore, that the probabilities in Figure 2 appear to be trending toward 1 and the values at the bottom of Table 10 indicate convergence to 1 in all cases.
@@
The results under all spatial models are shown in Table 11. The table provides four different probabilities for each model: the probability an election demonstrates a runner-up (RU) or a plurality loser (PL) reinforcement paradox (i.e., the probability an election demonstrates a reinforcement paradox), the probability an election demonstrates a runner-up paradox (RU), the probability an election demonstrates a plurality loser paradox (PL), and the probability an election demonstrates a double paradox (PL and RU). The first row gives unconditioned probabilities; the second row gives probabilities conditioned on the absence of a majority candidate for 1D spatial. The middle (respectively last) two rows give the corresponding probabilities under 1D Bimodal (respectively 1D Bimodal Weighted).
@@
If the distribution of voters is symmetric, $Z_{1}$ and $Z_{2}$ are likely to be similarly-sized, and so the rightmost inequality will be satisfied whenever $Y>X$. Thus an RU paradox will occur often when the IRV and plurality winners are distinct.
@@
For the PL paradox to occur, it is easily seen that $C$ must be the centrist candidate. In this case, $C$ is again likely to be on the upward slope of the larger peak in the 1D Bimodal Weighted model (and hence $Z$ is likely to be larger) than in the 1D Bimdodal model. So the probability of the PL paradox occurring is greater in the nonsymmetric case than the symmetric.
@@ 7.1. Theoretical Versus Empirical Results
What none of the models capture is that in actual elections, it is extremely rare to observe a plurality loser paradox without also observing a runner-up paradox. Under spatial models this difference is especially pronounced, as under those models if a plurality loser paradox is exhibited then we almost never also observe a runner-up paradox. This is generally true across all models, even though the RU paradox is roughly three times as likely to occur as the PL paradox.
@@ 8. Conclusion
Both theoretical and empirical analysis suggests that the reinforcement paradox is significantly more likely to occur under IRV with 3 candidates than other paradoxes. This is partly because, as noted above, there are no restrictions on how the ballots are partitioned. It is less likely, for instance, that a candidate will be the “victim” of a reinforcement paradox through winning two similarly-sized precincts despite losing when the ballots are combined. Future work could examine how the probability of a reinforcement paradox changes when the partition is restricted to sub-elections of approximately equal size, or to larger number of equally-sized sub-elections. Finally, this work could be extended to more than three candidates by determining conditions on the parameters in an $n$-candidate preference profile under which the paradox occurs.

### v2

@@ 5.2. Results
Throughout this section and the next, in our analysis we identify four probabilities: the probability that a runner-up (RU) paradox occurs; the probability that a plurality loser (PL) paradox occurs; the probability that at least one of these (RU or PL) paradoxes occurs; and the probability that both (RU and PL) paradoxes occur.
@@
Table 12 also displays information for three large $\alpha$ values which are not shown in Figure 2. Recall that as $\alpha\rightarrow\infty$ the distribution resembles the IC model, in which each voter independently chooses one of the six possible rankings at random. For comparison, we also conducted simulations using the IC model with $V=1001$, which returned an estimated probability of 1.0 for the runner-up, plurality, or both (RU, PL, and RU and PL) paradoxes occurring. Intuitively, this is clear since when every voter chooses a ranking at random, $X_{1}\approx X_{2}\approx Y_{1}\approx Y_{2}\approx Z_{1}\approx Z_{2}$. Thus, in the limit, Condition (2) and the plurality loser inequalities of Proposition 1 will always be satisfied. It is unsurprising, therefore, that the probabilities in Figure 2 appear to be trending toward 1 and the values at the bottom of Table 12 indicate convergence to 1 in all cases.
@@
The results under all spatial models are shown in Table 13. The first row gives unconditioned probabilities; the second row gives probabilities conditioned on the absence of a majority candidate for 1D spatial. The middle (respectively last) two rows give the corresponding probabilities under 1D Bimodal (respectively 1D Bimodal Weighted).
@@
If the distribution of voters is symmetric, $Z_{1}$ and $Z_{2}$ are likely to be similarly-sized, and so the rightmost inequality will be satisfied whenever $Y>X$. Thus a runner-up (RU) paradox will occur often when the IRV and plurality winners are distinct.
@@
For the plurality loser (PL) paradox to occur, it is easily seen that $C$ must be the centrist candidate. In this case, $C$ is again likely to be on the upward slope of the larger peak in the 1D Bimodal Weighted model (and hence $Z$ is likely to be larger) than in the 1D Bimdodal model. So the probability of the plurality loser (PL) paradox occurring is greater in the nonsymmetric case than the symmetric.
@@ 7.1. Theoretical Versus Empirical Results
What none of the models capture is that in actual elections, it is extremely rare to observe a plurality loser paradox without also observing a runner-up paradox. Under spatial models this difference is especially pronounced, as under those models if a plurality loser paradox is exhibited then we almost never also observe a runner-up paradox. This is generally true across all models, even though the runner-up (RU) paradox is roughly three times as likely to occur as the plurality loser (PL) paradox.
@@ 8. Conclusion
Both theoretical and empirical analysis suggests that the reinforcement paradox is significantly more likely to occur under IRV with 3 candidates than other paradoxes. This is partly because, as noted above, there are no restrictions on how the ballots are partitioned. If we require the two sub-elections to be similarly-sized, for example, as they might be in two precincts, there will most likely be fewer instances of the paradox; the same is true if we restrict the sub-elections so that one sub-election is a small perturbation of the original election. Future work could examine how the probability of a reinforcement paradox changes under restrictions of this kind, or to instances involving a larger number of sub-elections. Finally, this work could be extended to more than three candidates by determining conditions on the parameters in an $n$-candidate preference profile under which the paradox occurs.